OpenStudy (anonymous):

Helpful steps, please, A 37 foot ladder is leaning against the side of a building. if the ladder makes an angle of 24 degrees 31 inches with the side of the building. how far is the bottom of the ladder from the base of the building? ( I got 15.35 ft would this be correct?)

OpenStudy (anonymous):

i have no idea what "24 degrees 31 inches" means

OpenStudy (anonymous):

Me either.. That's why I was confused.

OpenStudy (anonymous):

my only guess would be 24 degrees 31 minutes, but who knows

OpenStudy (anonymous):

my mistake yes thats what I meant to state

OpenStudy (anonymous):

What did you do to solve this problem...Division, Multiplication, Adding, Subtracting?

OpenStudy (anonymous):

you gotta convert this to a decimal before you can calculate it

OpenStudy (anonymous):

then \[\cos(\theta)=\frac{x}{37}\] solve for \(x\) via \[x=37\cos(\theta)\]

jimthompson5910 (jim_thompson5910):

I'm guessing they gave you this angle: \(\Large 24^{\circ} 31'\) If so, then yes, that means "24 degrees, 31 minutes" \[\Large \theta = 24 \text{ degrees}, 31 \text{ minutes}\] \[\Large \theta = 24 \text{ degrees}, \frac{31}{60} \text{ degrees}\] \[\Large \theta = 24 \text{ degrees}+\frac{31}{60} \text{ degrees}\] \[\Large \theta = (24 +\frac{31}{60}) \text{ degrees}\] \[\Large \theta = (\frac{1440}{60}+\frac{31}{60}) \text{ degrees}\] \[\Large \theta = \frac{1440+31}{60} \text{ degrees}\] \[\Large \theta = \frac{1471}{60} \text{ degrees}\] \[\Large \theta \approx 24.5167 \text{ degrees}\]

OpenStudy (anonymous):

I understand that converting it to a decimal is 24.51667 would my equation be 37/cos(24.51667) ?

jimthompson5910 (jim_thompson5910):

do you see how satellite solved for x?

OpenStudy (anonymous):

Just that x= 37cos(theta) is that meaning that its 37cos(24.51667)?

jimthompson5910 (jim_thompson5910):

correct, you use a calculator at this point

OpenStudy (anonymous):

this is the issue. the answers are a)16.65 ft b)21.05ft c)15.35ft d)5.98ft and I am getting 33.66ft

OpenStudy (anonymous):

are you sure we use cos instead of using sin?

jimthompson5910 (jim_thompson5910):

let me re-read

jimthompson5910 (jim_thompson5910):

"A 37 foot ladder is leaning against the side of a building. if the ladder makes an angle of 24 degrees 31 inches with the side of the building. how far is the bottom of the ladder from the base of the building?" let x = distance from bottom of ladder to wall so you should have something like this 

jimthompson5910 (jim_thompson5910):

yeah you use sine instead of cosine

OpenStudy (anonymous):

so then 15.35ft?

jimthompson5910 (jim_thompson5910):

approximately, yes

OpenStudy (anonymous):

thanks for your help

jimthompson5910 (jim_thompson5910):

np